OpenStudy (anonymous):

Is the slope of 2x + 3 = 0 ... -3/2?

OpenStudy (anonymous):

this is not a function, hence there is no slope

OpenStudy (anonymous):

why is it not a function?

OpenStudy (anonymous):

function tend to be in form y(x) and f(x)

OpenStudy (anonymous):

in the given form 2x + 3 = 0 , the is an equation, I think you mean y = 2x + 3 where the slope id the coefficent of x since it;s in themx+b form

OpenStudy (anonymous):

but if you mean the equation 2x + 3 = 0 then the solution , that is the value of x will be as you said x=-3/2

OpenStudy (anonymous):

Yes the equation is 2x + 3 = 0. Yay I got it right!

OpenStudy (anonymous):

that's great